OpenStudy (ashy98):

The function below shows the number of students of a school who enrolled for ballet classes. Let f(x) represent the total number of students who enrolled for the classes after x years: f(x) = 12(1.06)x The average rate of change in the number of students who enrolled for ballet classes from the third to the fifth year is ________students per year. Round your answer to the nearest whole number.

OpenStudy (michele_laino):

please keep in mind that the average rate r is given by the subsequent formula: \[r = \frac{{f\left( 5 \right) - f\left( 3 \right)}}{{5 - 3}}\]

OpenStudy (michele_laino):

where f is your function

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

do you get how to do this?

OpenStudy (ashy98):

not really @misty1212

OpenStudy (misty1212):

lets use a calculator ok?

OpenStudy (ashy98):

okay(:

OpenStudy (ashy98):

-f(5)+5f(3)+r+3=0? so it would be 0?

OpenStudy (michele_laino):

no, I'm sorry I think that r is not zero

OpenStudy (ashy98):

f(x) = 12(1.06)^x f(x) = 12(1.06)^3 f(x) = 12(1.191016) f(x) = 14.292192 f(x) = 12(1.06)^x f(x) = 12(1.06)^5 f(x) = 12(1.33822558) f(x) = 16.058707 Average Rate of Change: m = y2 - y1/x2 - x1 Points: (3,14.292192) and (5,16.058707) m = 16.058707 - 14.292192/5 - 3 m = 1.766515/ 2 m = 0.8832575 So the average rate of change in the number of students who enrolled for ballet classes from the third to the fifth year is approximately 1 student per year.